OpenStudy (anonymous):

3x + 8y = 9 x = 22 - 9y What is the ordered pair?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

3x + 8y = 9 x +9y = 22 (-3) 3x +8y = 9 -3x -27y = -66 -------------- -19y = -57 y = 3 ............................. x = 22 - 9(3) x = 22 - 27 x = -5 (-5,3) is the ordered pair